So how would you calculate the force of F needed to keep the engine stationary by first resolving the forces into their horizontal and vertical components.


1
Expert's answer
2022-03-14T13:06:24-0400

Answer

If a force is at "\theta" angle from +x axis then it's horizontal and vertical components are

"F_H=F\cos\theta"

"F_v=F\sin\theta"